# IMAgent: Training Multi-Image Vision Agents via End2End Reinforcement Learning

### v2

@@ 4.2 Main Results
In Table 1 of the experimental results, we reveal a noteworthy phenomenon that, for base Qwen2.5-VL-7B, on V* and HR-Bench, directly generated answer (row 6) is superior to the step-by-step generated ones (row 7), especially on the higher resolution tasks(HR-Bench 8K).
On our MIFG-QA, where the input images have relatively high total resolution and abundant fine-grained information, we observe a similar phenomenon. We speculate that this phenomenon stems from the issue revealed in Figure 1.
While existing methods achieve significant improvements in single-image scenarios through external visual tools, they largely overlook multi-image tasks—either lacking native support or failing to evaluate in such settings. To benchmark these vision agents on our proposed MIFG-QA-val, we modify prompts for models incompatible with multi-image operations, forcing them to output image indices during tool call.
IMAgent leverages visual reflection and confirmation mechanisms to autonomously prioritize critical images/regions, surpassing state-of-the-art workflow-based methods [wu2024vstar, li2025dyfo, shen2025zoomeye] and training-based approaches [zhang2025chain, zheng2025deepeyes, su2025pixel] by significant margins. Crucially, this gain requires neither complex workflow design nor elaborate cold-start SFT, instead unlocking dynamic visual capabilities through end-to-end reinforcement learning.
@@ 4.3 Ablation Study
We further investigate the impacts of visual tools, training data, and trajectory masking on the model’s dynamic visual capabilities, with the results presented in Table 2.
A comparison between Row 3 and Row 5 demonstrates that when equipped with the visual reflection tool, the model’s overall performance (on the V*, MME, and MIFG-QA datasets) is significantly improved, despite a potential decrease on HR-Bench caused by the excessively high proportion of image tokens resulting from the reuse of high-resolution images. Explicit visual reflection enables the model to re-examine images, thereby achieving more precise alignment between questions and image regions. For multi-image scenarios, the reflection tool assists the model in identifying images relevant to the query.
Experiments in Row 2 and Row 5 show that incorporating multi-image data into the training process helps comprehensively enhance model performance. Multi-image data involves more image elements and comparisons between images, which effectively expands the diversity of tasks the model learns and encourages the model to explore richer tool use strategies. A comparison between RoW 4 and Row 5 indicates that the model’s multi-image processing capability drops sharply if trajectory masking is not applied during training. Multi-image tasks typically require more visual operations and interaction rounds. The lack of masking for invalid trajectories inhibits long-range interactions, thus limiting the model’s exploration of tool-use strategies for multi-image problems during training.
@@
5 Key Findings
5.1 Enabling Attention Reallocation via Visual Tools
Figure 5: Attention maps of the model using the visual confirmation tool. To the left of the blue dashed line are the query image pair, where the thumbnail in Text Box indicates the cropped region. To the right of the dashed line are the attention maps corresponding to one token each when the model outputs four bounding boxes.
@@ 5.1 Enabling Attention Reallocation via Visual Tools
Figure 6: Attention distribution across model layers. The first row shows the layer-wise attention distribution of all output tokens to the three token categories. The second row presents the layer-wise attention distribution of the output tokens where the model’s attention to images reaches its peak.
@@
We further analyze the attention distribution across all layers of the model during the entire reasoning process and at the moments of peak image attention, as shown in Figure 6. We found that higher image attention typically appears in the middle and later layers of the model, and IMAgent also exhibits significantly higher image attention in these layers. This phenomenon is more pronounced in the comparison of attention distribution at the token level: IMAgent demonstrates far greater visual attention than Qwen2.5-VL across multiple intermediate layers.
5.2 The Importance of Tool Incentives for Zero RL
Our experiments demonstrate that providing tool incentives for agents under the Zero RL paradigm is critical. Figure 7 illustrates the impact of tool gains on the agent’s ability to master tool invocation. When the tool incentive is set to 0, although the system prompt encourages the model to explore visual tools initially, this exploration gradually diminishes to 0 as the number of training steps increases—with the model preferring to learn direct reasoning strategies instead. This phenomenon may arise from three factors: the base model, training characteristics, and training data. Specifically, instruction-tuned models such as Qwen2.5-VL-Instruct lack interactive data related to tool use in their training corpora, leading to poor tool interaction capabilities. Additionally, exploratory tool use by the model in the early training stages often results in inaccurate or inefficient tool calls; the subsequent format penalties and feedback from failed tool invocations inhibit further tool exploration. Furthermore, certain samples in the training data do not strictly require visual processing to generate correct responses, enabling the model to bypass tool use and directly obtain answers.
Figure 7: Comparison of tools with and without trajectory masks.
To address this, we design an accuracy-driven tool use gain incentive for IMAgent. This incentive encourages the model to learn strategies that utilize tools to achieve correct responses while avoiding meaningless tool use. After introducing the tool gain, the average number of tool uses per step by IMAgent gradually increases and reaches a peak in the early training phase—indicating that the tool gain effectively motivates the model to explore tool use initially. In the middle training phase, the average number of tool uses per step gradually decreases, reflecting that the model has adapted to completing visual reasoning tasks with tools and begins to shift toward more efficient tool use strategies. In the late training phase, the model masters several reasoning strategies integrated with tool use, and the average number of tool uses per step exhibits a stable trend.
5.3 Zero RL inspires dynamic visual reasoning
Although IMAgent needs no SFT cold-start for memorizing specific reasoning pattern, end2end RL post-training still stimulates its dynamic visual reasoning ability. As shown in Figure 8, the model exhibits a usage pattern that prioritizes confirmation tools, supplemented by reflection tools, giving rise to diverse tool-use strategies exemplified by direct or exploratory visual confirmation, and cross-image comparative or reflective confirmation, among others.
Figure 8: Demonstration of some typical tool use strategies.
For single-image questions, when the question involves a moderately sized image region, IMAgent tends to adopt direct visual confirmation. It provides an answer after invoking the confirmation tool once to obtain feedback. If the target region is small or involves multiple comparative regions, the model tends to explore the target region or perform confirmation one by one to acquire detailed information. In such cases, the model may also conduct summary reflection after confirmation to verify the answer. For multi-image questions, the frequency of reflection tool use is higher than that in single-image scenarios. After performing attention-focused confirmation operations, the model may forget details of other images. At this point, the model will select the reflection tool to reallocate visual attention for image information, and then choose the confirmation tool again or directly answer the question. More detailed examples of reasoning are shown in the appendix.

### v3

@@ 4.2 Main Results
In Table 1 of the experimental results, we reveal a noteworthy phenomenon that, for base Qwen2.5-VL-7B, on V* and HR-Bench, directly generated answer is superior to the step-by-step generated ones, especially on the higher resolution tasks(HR-Bench 8K). On our MIFG-QA, where the input images have relatively high total resolution and abundant fine-grained information, we observe a similar phenomenon. We speculate that this phenomenon stems from the issue revealed in Figure 2.
While existing methods achieve significant improvements in single-image scenarios through external visual tools, they largely overlook multi-image tasks—either lacking native support or failing to evaluate in such settings. To benchmark these vision agents on multi-image data, we modify prompts for models incompatible with multi-image operations, forcing them to output image indices during tool call. However, some models, such as the DeepeyesV2 configuration code tool, are fundamentally unable to handle multiple image inputs.
Table 2: Average Tool Call Count by Benchmark. Values in superscript indicate (zoom in ratio / reuse ratio)
V* HR-Bench MME MIFGQA
1.00 (1.00 / 0.00) 1.05 (0.97 / 0.03) 1.12 (0.93 / 0.07) 1.70 (0.70 / 0.30)
Table 3: Ablation experiments: “text RL” means performing RL training on plain-text CoT without any tools,“w/o MIFG-QA” refers to adopting only single-image training data, “w/o lookback” denotes not using visual reflection tools, and “w/o Traj mask” signifies not masking invalid tracks.
Components V* HR-Bench MME MIFG-QA BLINK Mantis M4Bench
direct relative Avg 4K 8K nature poster Avg
Baseline 79.13 73.68 76.96 70.62 66.75 57.3 36.94 44.34 41.98 55.5 75.6 45.2
Baseline+text RL 82.61 78.95 81.15 72.75 68.75 63.57 45.81 42.26 43.40 55.2 77.4 48.9
IMAgent-7B w/o Tool gain 86.08 82.89 84.81 72.37 71.0 63.14 46.80 44.34 45.34 55.1 78.3 47.6
IMAgent-7B w/o MIFG-QA 87.83 80.26 84.81 72.75 70.5 64.02 39.90 40.18 40.09 56.1 77.0 53.5
IMAgent-7B w/o lookback 86.07 76.32 82.20 74.0 71.25 62.24 48.25 47.98 48.07 56.0 78.3 53.9
IMAgent-7B w/o Traj mask 86.07 81.59 84.29 73.5 69.75 64.03 46.31 46.18 46.23 56.7 78.3 54.5
IMAgent-7B 88.70 88.16 88.48 74.0 71.5 64.90 50.24 48.73 49.21 56.4 78.8 55.6
IMAgent Utilizes visual reflection and confirmation mechanisms to autonomously think about key images/regions, achieving state-of-the-art performance on single-image and multi-image tasks, and showing significant improvements on fine-grained tasks (V*, HR-Bench, MIFG-QA, M4Bench, etc.). In particular, we present tool usage statistics for the evaluation of IMAgent in the Table 2. For multi-image data, the proportion of visual reflection tool use increases significantly, indicating that the model tends to review images to avoid forgetting, thereby achieving more accurate reasoning in multi-image scenarios. Notably, IMAgent acquires dynamic visual reasoning capabilities and exhibits diverse tool-usage strategies through reinforcement learning alone. Representative cases include direct or exploratory visual confirmation, cross-image comparison, and cross-image reflective confirmation. More detailed reasoning examples are provided in the supplementary materials.
@@ 4.3 Ablation Study
We present the training curves of the average tool use counts for models with and without tool gain, as shown in Figure 6. When the tool gain is set to zero, the model’s exploratory utilization of tools gradually decays to zero as training steps increase. This phenomenon is common in non-native agentic models such as Qwen2.5-VL-Instruct (while not in Qwen3-VL). The pre-training corpus of these models lacks interactive data related to tool use, resulting in weak tool interaction capabilities. After introducing the accuracy-based tool use gain, the model is incentivized to learn strategies that leverage tools to derive correct answers while avoiding meaningless tool use. As training proceeds, the average number of tool use instances per step for IMAgent increases gradually and peaks in the early training stage—demonstrating that the tool gain effectively encourages the model to explore tool use during the initial phase. In the mid-training stage, the average per-step tool use count decreases steadily, indicating that the model has adapted to completing visual reasoning tasks with tool assistance and evolved toward more efficient tool use strategies. In the late training stage, the model masters diverse reasoning strategies integrated with tool use, and the average per-step tool use count tends to stabilize.
Figure 6: Comparison of tools with and without tool gain.
We further investigate the effects of visual tools, training data, and trajectory-level mask on the ability of Qwen2.5-VL-7B (which lacks native tool-usage capability) to learn dynamic visual skills, so as to verify the robustness of each component. The results are presented in Table 3.
A comparison between Row 5 and Row 7 demonstrates that when equipped with the visual reflection tool, the model’s overall performance (on the V*-relative, MME, MIFG-QA and M4Bench datasets) is significantly improved. Explicit visual reflection enables the model to re-examine images, thereby achieving more precise alignment between questions and image regions. For multi-image scenarios, the reflection tool assists the model in identifying images relevant to the query.
Experiments in Row 4 and Row 7 show that incorporating multi-image data into the training process helps comprehensively enhance model performance. Multi-image data involves more image elements and comparisons between images, which effectively expands the diversity of tasks the model learns and encourages the model to explore richer tool use strategies. A comparison between RoW 6 and Row 7 indicates that the model’s multi-image processing capability drops sharply if trajectory-level mask is not applied during training. Multi-image tasks typically require more visual operations and interaction rounds. The lack of masking for invalid trajectories inhibits long-range interactions, thus limiting the model’s exploration of tool-use strategies for multi-image problems during training.
@@
5 Dynamic Visual Attention in Reasoning
5.1 Visual Tools Enable Attention Reallocation
Figure 7: Attention maps for the visual confirmation tool. Left of the dashed line lies the query image pair; the text box thumbnail marks the cropped region. Right of the line are attention maps for one token each when the model outputs four bounding boxes.
@@ 5.1 Visual Tools Enable Attention Reallocation
Figure 8: Attention distribution across model layers.
@@
We further analyze the attention distribution across all layers of the model, as shown in Figure 8. The first row shows the layer-wise attention distribution of all output tokens to the three token categories. The second row presents the layer-wise attention distribution of the output tokens where the model’s attention to images reaches its peak. We found that higher image attention typically appears in the middle and later layers of the model, and IMAgent also exhibits significantly higher image attention in these layers. This phenomenon is more pronounced in the comparison of attention distribution at the token level: IMAgent demonstrates far greater visual attention than Qwen2.5-VL across multiple intermediate layers.
5.2 Visual Tools Drive Attention Focus on Task-Relevant Regions
Figure 9: Results of average attention ratio across visual regions. ROI denotes tool-call bounding boxes region; GT indicates original answer-related dataset regions.
[…]
Figure 9 (a) shows that in the tool use setting, $F(\text{ROI})$ during the <tool_call> phase is substantially higher than during the <think> phase. This indicates a natural decoupling between perception and reasoning: the model leverages tool calls for targeted visual perception while using the think phase for language-level reasoning.
Figure 9 (b) reveals that the elevated $F(\text{GT})$ during <tool_call> confirms that the ROI regions queried by tool calls are highly correlated with the answer-relevant areas. This suggests that the performance gains of IMAgent stem from more precise visual localization of task-relevant regions. Regarding the <Avg> score in (b), the baseline Qwen2.5-VL achieves $F(\text{GT})=10$ without tool use. Equipping the same model with tools raises this to $19.0$, while IMAgent further improves it to $26.9$. This demonstrates that explicit tool calling can strengthen a model’s visual attention toward answer-relevant regions, and that this effect can be further amplified through training.
Overall, IMAgent consistently achieves higher attention ratios across both ROI and GT regions, with the most pronounced improvement in <tool_call> tokens. These results confirm that IMAgent actively redirects visual attention to task-relevant regions throughout inference, rather than relying on a fixed first impression of the image.